40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 7 of 339  Not logged in ELOG logo
ID Date Author Type Category Subject
  16783   Mon Apr 18 14:52:47 2022 Ian MacMillanSummarySEISeismic Study of Buildings and Caltech Campus

[Ian, JC]

I want to take measurements of seismic noise at different places on Caltech's campus and in different buildings. I will try to use the accelerometer in my phone for this but first I must calibrate it (Against the 40m accelerometers). 

I placed my iPhone 11 pro next to the seismometers at the 40m MC as seen in Attachment 1.

The calibration from the instrument was done using cts/rthz * 1V/16384cts * 1/ampgain * g/10V * 10m/s^2/g. The ampgain for all was 100.

Next, I took 100 seconds of data on both the iPhone and the three orthogonal Wilcoxon accelerometers.

The ASD for both of the total acceleration is shown in Attachment 2

The ASD for the individual directions acceleration is shown in Attachment 3

The coherence between the individual directions acceleration and the 40m's individual directions is shown in Attachment 4. For this calculation, the 40m data were downsampled to roughly match the phone's sample rate. This coherence is not very good. It should be higher. Because the phone and 40m sensors were picking up the same data as the phone. Because of this I also looked at the coherence between the individual 40m sensors.

In Attachment 5 I look at the coherence between the individual 40m sensors. This should give me a good idea of whether this is some other issue giving me mow coherence. This plot shows that the coherence between the individual 40m sensors is much better than between the phone and the 40m sensors.

Now I wanted to see what kind of data the iPhone could get from real-world tests. I placed it in a number of locations described below and plotted their ASDs in Attachment 6. The locations are thus:

Identifier  Location Notes
QIL QIL Lab in the Sub-basement of west bridge In sub-basement not much activity when taking measurements.
WBSH West bridge sub-basement hallway on floor in hallway no activity around
WB1H West Bridge 1st floor Hall placed on the floor near pillar near stairs to LIGO offices on the ground floor of west bridge
40m desk on my desk at the 40m placed on the desk while people were walking around and I had my feet on the desk. should be noisy

Notice how at the low end the amplitudes follow the relative amplitudes I would expect. QIL and WBSH are the lowest then WB1H is noisier and 40m desk is the noisiest. However, this is only true up until about 0.5 Hz then they all overlap. Since I would expect the 40m desk should be much noisier at all frequencies I suspect that the phone accelerometer is not suitable for measurements higher than 0.5 Hz.

Possible Problems:

One possible problem with my measurement is that my phone was in a leather case. this may have damped out higher frequencies. Also, my phone was not weighed down or bolted to the floor. this stronger connection would make it better at detecting higher frequencies. I could repeat the experiment with no case and a weight on top of my phone.

What's next:

Since I don't think the phone can give me accurate data above 0.5Hz for quiet environments. It may not be suitable for this task. It would seem that the right instrument is the Wilcoxon 731A but it requires an amplifier that I can't track down.

 

I included all the data and code in the zip file in attachment 7

 

Attachment 1: IMG_0513.jpg
IMG_0513.jpg
Attachment 2: tot_acc_cal.pdf
tot_acc_cal.pdf
Attachment 3: indiv_acc_cal.pdf
indiv_acc_cal.pdf
Attachment 4: dec_Coherence.pdf
dec_Coherence.pdf
Attachment 5: 40m_self_Coherence.pdf
40m_self_Coherence.pdf
Attachment 6: tot_acc_testsites.pdf
tot_acc_testsites.pdf
Attachment 7: Calibration.zip
  16782   Fri Apr 15 11:59:16 2022 YehonathanUpdateIOOIMC completely misaligned

Came this morning, opened the PSL and there was not even a beam on the MC REFL.

Looking at the big monitor it seems like the WFS signals went through the roof during the "auto-alignment" night session.

I restored the MC alignment from before the misalignment happen and wait for the SUS to damp. Once the RMS values went below 200 I enabled the watchdog and the coil outputs.

I opened the PSL shutter and the IMC locked immediately. I turned on the WFS servo and the MC REFL DC went down to 0.3. I run the WFS relief script.

  16781   Thu Apr 14 18:39:13 2022 PacoSummaryBHDPart IIa of BHR upgrade - POY11 debugging

[Paco]

In reply to Koji's questions;

Q1: What is the product number of the broken Kepco?

A1: Kepco JQE 0-25 V 4A (1/4 rack mountable 100 W linear power supply)

Q2: Do you feel the burning smell on this broken kepco?

A2: Not very clearly. It just smells like generic broken electronic, but the fan was already long gone by the time we detected its problem.

Q3: How much current does the +15 VDC line draw from the Sorensen?

A3: The Sorensen current monitor reads 6.3 Amps

Q4: Is there a linear power supply in the market that can handle this much current with some margin?

A4: Probably... but we would have to look around.

Q5: Do we really need a linear power supply there?

A5: Good question, I guess anything that is not contaminating the RF electronics with HF noise (e.g. switching PS) can work?

Q6: Is the same fan problem happening on other Kepcos?

A6: Other fans are on, but at least one or two have the "ill" sound... It may be worthwile to give them maintenance if we can.

  16780   Thu Apr 14 18:34:51 2022 PacoSummaryBHDITMY Oplev reinstalled (Re: 2 in oplev mirrors incompatible with LMR2V)

[Paco, Yehonathan]

We installed ITMYOL1 and ITMYOL2 on the ITMY chamber. We aligned the ITMY OpLev beam and closed the loop successfully, we then had a second round of YARM aligment, where we brought the Y peak transmission up from 0.04 counts to 0.09 counts (up by a factor of two). We still couldn't close the YARM loop but we have a better alignment.

  16779   Thu Apr 14 11:52:57 2022 yehonathanSummaryBHDPart IIa of BHR upgrade - POY11 debugging

{JC, Paco, Yehonathan, Ian}

POY lens was moved to infront of the POY steering mirror to make the POU beam focused on the POY11 RFPD. We measured the DC output with an oscilloscope and optimized it with the steering mirrors. We get ~ 16.5mV.

The new lens position blocked the BS OpLev ingoing beam, so we repositioned the OpLev mirrors to make the beam path not hit the lens.

We went to the control room to observe the PDH signal. We observed a series of PDF osscillation and then the signal died infront of our eyes! There is just noise.

We go and check the +/-15V powering the RFPD and we find that the V- is ~ 14V which is good but the V+ was ~ 2.7V which is not.

We went to the PD interface and measure the POY11 output oltages using a breakout board and got the same result.

The PD interface was taken out for inspection. All the OP27 on channel 3 were replaced with new ICs (without need turns out)...


The PD interface card turned out to be OK. What happened is that one of the Kepcos in the RF rack died because its fan crumbled as seen in Attachment #2 (could this be the source of burning smell?). In response, the rack was drawing from the other Kepco (connected in parallel) way too much current  (4A) and the current limiter dropped its voltage from 15V to 2.7V.

The Kepco pair was removed and replaced with a single Sorensen. The POY PDH signal was restored (see attachment).

Attachment 1: Screenshot_2022-04-14_15-53-39.png
Screenshot_2022-04-14_15-53-39.png
Attachment 2: PXL_20220414_220616875.jpg
PXL_20220414_220616875.jpg
  16778   Thu Apr 14 10:18:35 2022 PacoSummaryBHD2 in oplev mirrors incompatible with LMR2V

[Paco, JC]

We realized the 2 in oplev mirrors (Thorlabs BB2-E02) for ITMYOL, SRMOL, and BSOL, are 0.47 in thick, while the LMR2V fixed mount is 0.46 in deep, even without taking the retaining ring into account. After a brief exchange with Koji, and Ian, we decided to glue the mirrors onto the mounts using Torr Seal (a vac compatible epoxy). They are curating in the clean room and should be ready to install in about 2 hours.

  16777   Thu Apr 14 09:04:30 2022 JordanUpdateVACRGA Volume RGA Scans

Prior to venting the RGA volume on Tuesday (4/12/2022) I took an RGA scan of the volume to be vented (RGA+TP1 volume+Manual Gate Valve) to see if there was a difference after replacing the manual gate valve. Attached is the plot from 4/12/22, and an overlay plot to complare 4/12/22 to 12/10/2021, when the same volume was scanned with the old (defective) manual gate valve.

There is a significant drop in the ratio O2 compared the the nitrogen peak and reduced Argon (AMU 40) which indicates there is no longer a large air leak.

12/10/21 N2/O2 ratio ~ 4 (Air 78%N2 / 21%O2)

4/12/22 N2/O2 ratio ~ 10      

There is one significant (above noise level) peak above AMU 46, which is at AMU 58. This could possibly be acetone (AMU 43 and 58) but overall the new RGA Volume scans look significantly better after the manual gate valve replacement. Well done!

Attachment 1: 40mRGA_Overlay.pdf
40mRGA_Overlay.pdf
Attachment 2: RGAVolume_4_12_22.PNG
RGAVolume_4_12_22.PNG
  16776   Wed Apr 13 18:55:54 2022 KojiUpdateCamerasCamera Battery Test

I believe that the Nikon has an exposure problem and that's why we bought the Canon.

 

  16775   Wed Apr 13 16:23:54 2022 Ian MacMillanUpdateGeneralSmell in 40m

[Ian, Paco, JC]

There is a strange smell in the 40m. It smells like a chemically burning smell maybe like a shorted component. I went around with the IR camera to see if anything was unusually hot but I didn't see anything. The smell seems to be concentrated at the vertex and down the y-arm

  16774   Wed Apr 13 15:57:25 2022 Ian MacMillanUpdateCamerasCamera Battery Test

Tested the Nikon batteries for the camera. they are supposed to be 7V batteries but they don't hold a charge. I confirmed this with multi-meter after charging for days. Ordered new ones Nikon EN-EL9

  16773   Wed Apr 13 12:50:07 2022 TegaUpdateVACNew Pressure Gauge Install/Pump Spool Vent

[Jordan, JC, Tega]

We have installed all the FRGs and updated the VAC medm screens to display their sensor readings. The replacement map is CC# -> FRG#, where # in [1..4] and PRP1 -> FRG5. We now need to clean up the C1VAC python code so that it is not overloaded with non-function gauges (CC1,CC2,CC3,CC4,PRP1). Also, need to remove the connection cables for the old replaced gauges.

Attachment 1: VAC_FRG_UPGRADE.png
VAC_FRG_UPGRADE.png
  16772   Tue Apr 12 09:05:21 2022 JordanUpdateVACNew Pressure Gauge Install/Pump Spool Vent

Today, Tega and I would like to vent the pump spool an dinstall the new FRG-400 Agilent Pressure Gauges (per elog 15703). The attached picture shows the volume needed to be vented highlighted in red, and the gauges that need to be replaced/removed (purple dot next to the name).

The vent plan is as follows:

Open RV2

Open VM3

Open V7

Open V4

Shut down TP2

Install new gauges

Will add to post with updates post vent.

Attachment 1: Screenshot_2022-04-12_08-42-33.png
Screenshot_2022-04-12_08-42-33.png
Attachment 2: 81CB0936-1B19-4722-8A32-C3DC1D1FBC21.heic
Attachment 3: 2262ECEF-6200-4E95-8E32-C83CB9EB4F17.heic
Attachment 4: B5712B34-ECF6-43BE-BCB2-38CD775CF653.heic
Attachment 5: BEE14A07-976A-45C2-82A0-1774D377941E.heic
Attachment 6: 84557D6E-6AAE-47CF-A3AD-1DF329FEB550.heic
  16771   Mon Apr 11 21:44:14 2022 KojiSummaryBHDPart IIa of BHR upgrade - POY11 debugging

We took out the right most PD interface board D990543 and removed the 74LS04 chips. In fact two out of 4 were already replaced with enabling wires, however it seemed that one of the remaining two got failed.
These remaining two chips were removed and the enabling signals were connected to VCC (+5V). This operation made the status LED light not functional. (We didn't bother to fixed them by connecting them to the GND)

The new schematic diagram was attached here, and the corresponding DCC entry (https://dcc.ligo.org/D1900318-v1) was modified. Attachments #2-3 show the circuit after the changes, and the front view respectively.

Attachment 1: D990543A1.PDF
D990543A1.PDF D990543A1.PDF D990543A1.PDF
Attachment 2: PXL_20220412_023941730.jpg
PXL_20220412_023941730.jpg
Attachment 3: PXL_20220412_023950567.jpg
PXL_20220412_023950567.jpg
  16770   Mon Apr 11 21:13:21 2022 PacoSummaryBHDPart IIa of BHR upgrade - POY11 debugging

[Paco, Koji]

I asked Koji for some advice regarding closing the loop on YARM using POY11. A few things seemed off including

  • The YARM transmission; which was peaking at ~ 20 (typically, TRY is normalized so that under nominal input power conditions we see TRY in the range [0, 1])
  • The POY11 DCPD level was quite low; we expect a few tens of uW in this low power configuration where no more than 100 mW are going through IMC.

We looked at the POY11 RFPD first. We tried flashing an incandescent lamp in-situ and saw some weak response using an oscilloscope and the DC Out readout. We then used the OPHIR power meter and recorded ~ 1.2 uW of light incident on the POY11 RFPD... Initially, we suspected our beam was not filling the PD sensitive area (~ 2 mm diameter), but a quick estimate using the 200 mm focusing lens currently installed in the ITMY table gave us quite a generous margin of error... so we questioned the OPHIR measurement. We swapped the power meter and this time got ~ 18.4 uW, which is more in line with what we expected (phew).

Moving on, from the POY11 RFPD responsivity, and our ~ 20 uW of incident power, we expected on the order of a 20 mV of DC Output, but weren't really seeing this on the scope, so we decided to test the pins on the power of the RFPD. The DB15 cable not only supplies bipolar 15 VDC but also monitors several other test points such as Tsens, or DCout in the RFPD. We quickly noticed a weird signal on the ENAB testpoint, so we removed POY11 RFPD from the ITMY table and took it to the PD testbench. After redoing the soldering on the breakout board and RF amplifier (ZXX-500LN+), which we tested separately, we saw the expected behavior using a +- 15 VDC power supply... thus verifying that the RFPD and breakout board seemed to work ok. We turned our attention to the upstream DB15 connection, and after quickly checking the newly run cables, we ended up debugging the eurocrate PD interface. After attempting a simple power cycle and failing, we removed this card and looked at the schematic. It would seem that the logic enabling ICs (one or both) failed, thus preventing the card from enabling its outputs correctly.... We bypassed the logic by soldering pins 4,8,14 on U1, U20 and then checked the circuit in-situ, and we saw it worked fine again.

Now none of the status LEDs (which are driven by the logic IC portion) work on this card, but the card itself works fine at least for POY11.

We moved on, and installed the DB15 cables, checking the functionality at every step... Then we looked at the POY11_I_ERR signal and were happy to see nice pdh wavelets. We pushed forward a little bit more to try and lock YARM. First, we went to the Y end and centered the ETMY Oplev so as to register the position where the YARM is flashing... The ITMY Oplev is still not online. Then, we optimized the ETMY damping gains somewhat to try and make it less noisy, and finally, played with the LSC YARM loop gain to attempt locking. This last push was not as successful, but we have an idea of what next steps are needed to reduce the SUS noise, including

  • Install ITMY Oplevs, and close loop
  • Optimize ITMY damping gains

To be continued....

  16769   Mon Apr 11 11:00:30 2022 JancarloUpdateVACC1VAC Reboot and Nitrogen tanks

[Paco, JC, Ian, Jordan, Chub]

Checking in the morning, I walked over to the Nitrogen tanks to check the levels. Noticed one tank was empty, so I swapped it out. Chub came over to check the levels and to take note of how many tanks were left available for usage (None). Chub continued to put in a work order for a set of full Nitrogen tanks. We should be set on Nitrogen until Thursday this week (4/14/22).

As for C1VAC, this morning, Paco and I attempted to open the PSL shutter, but the interlock system was tripped so we didn't get any light into the IFO. We traced the issue down to C1VAC being unresponsive. We discussed this may have interlocked as a result of the Nitrogen tanks running out, but we do not believe this was the issue since we would have recieved an email. We tried troubleshooting as much as possible avoiding a reboot, but were unable to solve the issue. In response, we ran the idea of a reboot across Jordan and Ian, where everyone was in agreement, and fixed the system. Restarting c1vac seems to have closed V4, but this didn't cause any issues with the current state of the vacuum system.

After opening the PSL shutter again, we see the laser down the IFO, so we resume alignment work

  16768   Fri Apr 8 17:21:31 2022 Ian MacMillanSummaryBHDPart IIa of BHR upgrade - IR laser alignment on Yarm

[Ian, Paco, Tega]

Paco and I opened the ETMY and ITMY chamber to work on yesterdays efforts to lock the y arm. We temporarily in stalled a camera behind the ETMY to look at the transmission as we adjusted the ETMY and ITMY. We then moved on to setting up the POY. the beam was too large for the apature of the PD so we installed a lens in the beam path to decrease it.

Once that was installed we saw some flashing on the C1:LSC-POYCD_OUT channel. We also could see the flashing on the monitors in the control room. The flashing beam seemed to be in the middle of ITMY but was slightly to the right on the ETMY. From here we tried to walk the beam using PR3 and ETMY to move the beam to the center of the ETMY.

  16767   Fri Apr 8 16:03:58 2022 ranaUpdateElectronicsRFSoC 2x2 board -- setup for remote work & BALUN saga

Seems like it should be possible to just remove the transformer (aka as a BALUN ... BALanced, UNbalanced), or replace it with a lower frequency part. Its just a usual mini-circuits part. Maybe you can ask Chris Stoughton about this and ask Tommy to checkout some of the RFSoC user forums for how to go to DC.

Quote:
 

After fiddling around with the tone-generators and spectrum analyzer tools in loopback configuration (DAC --> ADC direct connection), we noticed that lower frequency (~ 1 MHz) signals were hardly making it out/back into the board... so we looked at some of the schematics found here and saw that both RF data converters (ADC & DAC) interfaces are AC coupled through a BALUN network in the 10 - 8000 MHz band (see Attachment #1). This is in principle not great news if we want to get this board ready for audio-band DSP.

We decided that while Tommy works on measuring TFs for SHP-200 all the way up to ~ 2 GHz (which is possible with the board as is) I will design and put together an analog modulation/demodulation frontend so we can upconvert all our "slow" signals < 1MHz for fast, wideband DSP. and demodulate them back into the audio band. The BALUN network is pictured in Attachment #2 on the board, I'm afraid it's not very simple to bypass without damaging the PCB or causing some other unwanted effect on the high-speed DSP.

 

  16766   Thu Apr 7 21:15:04 2022 YehonathanUpdateBHDInitial BHD modeling: AS - LO mode matching

Ok, it turns out these optics were purchased on purpose, as this elog shows. Jon considered building a mode-matching telescope with stock optics as an initial step before purchasing the custom optics (referred to as "design" optics in my elog).

I dialed in the new distances between the optics into the .kat file as described in this elog and pushed the changes to the repo. With the new distances, I got mode-matching of 87% for the full IFO and 89% for FPMI so there's probably no need to worry as the mode-matching with these optics was already designed.

Quote:

I begin modeling the initial BHD setup using Finesse. I started with copying C1_w_BHD.kat from the 40m/bhd repo and making changes to reflect the current BHD setup:

1. OMCs were removed.

2. Only 1 PD per BHD port was left.

3. Transmission of PR2 was changed to 2.2%. The PRG was calculated to be ~15.5.

4. Actual RoCs of new optics were dialed in (Yesterday me and Paco went into the cleanroom to measure the RoCs and they seem to match the datasheets).

Here's a table comparing the old (design?) RoCs with the new RoCs:

  New RoC Old RoC
LO1 5m 6m
LO2 inf inf
LO3 500mm 750mm
LO4 150mm -450mm
AS1 2m 2.8m
AS2 inf inf
AS3 200mm -2m
AS4 750mm 600mm
PR2 2000m -700m
PR3 1000m 1000m
SR3 1000m -700m

 

The changes looked quite alarming, especially for LO4 and AS3, so I wrote a script to calculate the mode matching between the LO and AS beams called AS_LO_ModeMatching.ipynb and pushed it into the repo. In the notebook a bright AS beam is created by creating a small asymmetry between the arms of ~ 0.003 degrees (~10pm). Amplitude detectors were put at the input ports of the BHD BS to calculate the fields in the AS and LO beams. In particular TEM00, TEM02 and TEM20 were measured for each beam.

The calculation shows that with the old RoCs the mode matching between the LO and AS beams is 99% while for the new RoCs it is ~ 50%.

 

  16765   Thu Apr 7 20:41:15 2022 TommyUpdateElectronicsRFSoC 2x2 Board -- Gain Plotter

In this file (under Tommy), we have a notebook which runs through a spectrum of frequencies and determines the gain response of the attached filter. Below we have the output of a high pass filter. We use IQ demodulation to change IQ componets to DC. Then using a butterworth filter, we read out the DC components and determine the gain's magnitude and phase. However, the phase seems very noisy. This is because the oscillators in the different tiles are independent and a random phase is introduced by changing the mixer frequency in individual tiles. To resolve this we need Multi Tile Synchronization or "MTS". 

Original Pynq Support Forum Query: https://discuss.pynq.io/t/rfsoc-2x2-phase-measurement/3892

We also have the code to fit a resposne function using IIRregular, but this is not as useful without proper phase data.

Attachment 1: Screen_Shot_2022-04-07_at_8.40.23_PM.png
Screen_Shot_2022-04-07_at_8.40.23_PM.png
Attachment 2: Unknown-3.png
Unknown-3.png
Attachment 3: Unknown-4.png
Unknown-4.png
  16764   Thu Apr 7 20:37:06 2022 TommyUpdateElectronicsRFSoC 2x2 Board -- Simple Tone Generator

In the "Tommy" sub folder, I created a new notebook called "SimpleToneGenerator". This tunes the DAC and ADC mixers to a single frequency and reads off the Time Series and Fourier components. We can alos easily check the demodulation scheme and implement butterworth filters to check their function.

  16763   Thu Apr 7 20:33:42 2022 TommyUpdateElectronicsRFSoC 2x2 board -- setup for remote work

To access the board remotely through the 40m lab ethernet port, use

ssh -N -L localhost:1137:localhost:9090 xilinx@<ip_address>

Then in the browser go to

localhost:1137/lab

Other SSH commands using different ports or without the -N -L seemed to fail to open Jupyter. This way has been successful thereafter.

Quote:

[Tommy, Paco]

Since last week I've worked with tommy on getting the RFSoC 2x2 board to get some TFs from simple minicircuits type filters. The first thing I did was set up the board (which is in the office area) for remote access. I hooked up the TCP/IP port to a wall ethernet socket (LIGO-04) and the caltech network assiggned some IP address to our box. I guess eventually we can put this behind the lab network for internal use only.

After fiddling around with the tone-generators and spectrum analyzer tools in loopback configuration (DAC --> ADC direct connection), we noticed that lower frequency (~ 1 MHz) signals were hardly making it out/back into the board... so we looked at some of the schematics found here and saw that both RF data converters (ADC & DAC) interfaces are AC coupled through a BALUN network in the 10 - 8000 MHz band (see Attachment #1). This is in principle not great news if we want to get this board ready for audio-band DSP.

We decided that while Tommy works on measuring TFs for SHP-200 all the way up to ~ 2 GHz (which is possible with the board as is) I will design and put together an analog modulation/demodulation frontend so we can upconvert all our "slow" signals < 1MHz for fast, wideband DSP. and demodulate them back into the audio band. The BALUN network is pictured in Attachment #2 on the board, I'm afraid it's not very simple to bypass without damaging the PCB or causing some other unwanted effect on the high-speed DSP.

 

  16762   Thu Apr 7 17:59:51 2022 AnchalSummaryBHDPart IIa of BHR upgrade - IR laser alignment on Yarm

[Anchal, Paco, JC, Tega]

Today we have aligned the Yarm cavity for IR, with verified 1.5 roundtrips. We also placed following optics in the BS table.

  • POXM1 (installed earlier, not eloged)
  • SRMOL2
  • POYM1
  • SRMOL1
  • POYM2

We also cleared the in-air BS table of all previous optics. JC and Tega setup HeNe laser for Oplevs roughly for now. Tega also transported the POY RFPD from ITMY in-air table to the BS in-air table. We aligned the POY path to the table, but we had to move ITMY after that to get 2nd roundtrip in the arm cavity, which misaligned our POY path again. POY path would need to be modified tomorrow.

 

  16761   Thu Apr 7 11:47:22 2022 YehonathanUpdateBHDInitial BHD modeling: AS - LO mode matching

I begin modeling the initial BHD setup using Finesse. I started with copying C1_w_BHD.kat from the 40m/bhd repo and making changes to reflect the current BHD setup:

1. OMCs were removed.

2. Only 1 PD per BHD port was left.

3. Transmission of PR2 was changed to 2.2%. The PRG was calculated to be ~15.5.

4. Actual RoCs of new optics were dialed in (Yesterday me and Paco went into the cleanroom to measure the RoCs and they seem to match the datasheets).

Here's a table comparing the old (design?) RoCs with the new RoCs:

  New RoC Old RoC
LO1 5m 6m
LO2 inf inf
LO3 500mm 750mm
LO4 150mm -450mm
AS1 2m 2.8m
AS2 inf inf
AS3 200mm -2m
AS4 750mm 600mm
PR2 2000m -700m
PR3 1000m 1000m
SR3 1000m -700m

 

The changes looked quite alarming, especially for LO4 and AS3, so I wrote a script to calculate the mode matching between the LO and AS beams called AS_LO_ModeMatching.ipynb and pushed it into the repo. In the notebook a bright AS beam is created by creating a small asymmetry between the arms of ~ 0.003 degrees (~10pm). Amplitude detectors were put at the input ports of the BHD BS to calculate the fields in the AS and LO beams. In particular TEM00, TEM02 and TEM20 were measured for each beam.

The calculation shows that with the old RoCs the mode matching between the LO and AS beams is 99% while for the new RoCs it is ~ 50%.

  16760   Wed Apr 6 22:51:47 2022 KojiSummaryBHDPart IIa of BHR upgrade - IR laser alignment on Xarm

[Yuta Koji]

We took out the two coil driver units for PR3 and the incorrect arrangement of the output Rs were corrected. The boxes were returned to the rack.

In order to recover the alignment of the PR3 mirror, C1:SUS_PR3_SUSPOS_INMON / C1:SUS_PR3_SUSPIT_INMON / C1:SUS_PR3_SUSYAW_INMON were monitored. The previous values for them were {31150 / -31000 / -12800}. By moving the alignment sliders, the PIT and YAW values were adjusted to be {-31100 / -12700}. while this change made the POS value to be 52340.

The original gains for PR3 Pos/Pit/Yaw were {1, 0.52, 0.2}. They were supposed to be reduced to  {0.077, 0.04, 0.015}.
I ended up having the gains to be {0.15, 0.1, 0.05}. The side gain was also increased to 50.

----

Overall, the output R configuration for PR2/PR3 are summarized as follows. I'll update the DCC.

PR2 Coil Driver 1 (UL/LL/UR) / S2100616 / PCB S2100520 / R_OUT = (1.2K // 100) for CH1/2/3

PR2 Coil Driver 2 (LR/SD) / S2100617 / PCB S2100519 / R_OUT = (1.2K // 100) for CH3

PR3 Coil Driver 1 (UL/LL/UR) / S2100619 / PCB S2100516 / R_OUT = (1.2K // 100) for CH1/2/3

PR3 Coil Driver 2 (LR/SD) / S2100618 / PCB S2100518 / R_OUT = (1.2K // 100) for CH3

Attachment 1: PXL_20220406_213356237.MP.jpg
PXL_20220406_213356237.MP.jpg
Attachment 2: PXL_20220406_212936502.MP.jpg
PXL_20220406_212936502.MP.jpg
  16759   Wed Apr 6 12:03:51 2022 AnchalSummaryBHDPart IIa of BHR upgrade - IR laser alignment on Yarm

[Anchal, Paco]

After Koji reduced the output resistors on PR2/PR3 coil drivers, we got much better actuation range. We aligned TT1 TT2 again to get beam centered on PR2 and PR3. Then we used only PR2 and PR3 to do the input beam alignment to Y arm cavity. Using access in ETMY chamber, we aligned the input beam parallel to cavity axis. Slight changes were required in ETMY alignment offsets to get first roundtrip in same spot on ITMY. remaining alignment is finer and needs to be done with a help of a reflection photodiode and cameras in the control room. Immediate next step is to setup POY path for locking the Yarm with IR.

Side note: Because of the large PIT correction required in PR3, we found that our upper OSEMs were hitting totally bright limit and lower OSEMs were hitting totally dark limit on PR3. This also destablized our damping loops. We pushed the upper OSEMs slighlty and pulled back the lower OSEMs slightly to get the PD signal in half shadow region again. This worked and our damping loops are stable again. However, we think we should repeat free swing test in future to diagonalize the input matrix for new OSEM positions.

  16758   Wed Apr 6 01:20:48 2022 KojiSummaryBHDPart IIa of BHR upgrade - IR laser alignment on Xarm

PR2/PR3 Output R for fame OSEMS reduced from 1.2K to 1.2K//100Ohm

I put the R=100Ohm for PR3 with the functions of the units mistakenly swapped. This affects imbalanced actuation of PR3 right now as well as too strong SD

PR2 Coil Driver 1 (UL/LL/UR) / S2100616 / PCB S2100520 / R_OUT = (1.2K // 100) for CH1/2/3

PR2 Coil Driver 2 (LR/SD) / S2100617 / PCB S2100519 / R_OUT = (1.2K // 100) for CH3

PR3 Coil Driver 1 (UL/LL/UR) / S2100619 / PCB S2100516 / R_OUT = (1.2K // 100) for CH3 only

PR3 Coil Driver 2 (LR/SD) / S2100618 / PCB S2100518 / R_OUT = (1.2K // 100) for CH1/2/3

----

The output R was reduced from 1.2k to 1.2k//100 = 92 Ohm.

This means that the face coil gains were increased by a factor of 13.

The original gains for PR2 Pos/Pit/Yaw were {0.7, 0.3, 0.2}. To keep the same loop gain, the new gains were supposed to be {0.054, 0.023, 0.015}.
With the new gain, the oscillations were very slowly reduced. Therefore, I increased the gains to have the gain margin of 2. (i.e. increased the gains until I have the oscillation, and then made it half.)
The new values were {0.2, 0.1, 0.05}. The side gain was 20 and unchanged

For PR3 the same operation has been done.

The original gains for PR3 Pos/Pit/Yaw were {1, 0.52, 0.2}. They were supposed to be reduced to  {0.077, 0.04, 0.015}.
The gains were increased to {0.5, 0.1, 0.1}. The side gain was also increased from 1 to 5.

Attachment 1: PXL_20220406_060846581.jpg
PXL_20220406_060846581.jpg
Attachment 2: PXL_20220406_014205726.jpg
PXL_20220406_014205726.jpg
Attachment 3: PXL_20220406_060324812.jpg
PXL_20220406_060324812.jpg
Attachment 4: PXL_20220406_015259326.jpg
PXL_20220406_015259326.jpg
  16757   Tue Apr 5 18:15:06 2022 AnchalSummaryBHDPart IIa of BHR upgrade - IR laser alignment on Xarm

[Paco, Anchal, Ian, JC]

We attempted the alignment of IR beam into the arm cavities. We used PR2 and PR3 (moved manually as well as using cdsutils) and got the YAW aligned pretty good on both X and Y directions. PIT alignment however turned out to be much harder to align. PR2 PR3 didn't have much range, so we zeroed there offset and tried to use TT1, TT2, MMT1, and MMT2 to align the PIT but it would get clipped before reaching BS table if we were to correct for PIT misalignment happening downstream. We concluded that the issue is that one of the PR2, PR3 mirrors have too much PIT offset in equilibrium position. We have requested Koji to change the output resistors in the coil drivers of PR2 and PR3 so that we can correct for the PIT offset in them directly using the coils and reduce load on upstream optics. We have tweaked TT1, TT2, MMT1, and MMT2 positions today, so we do not have the previous reference anymore.

  16756   Mon Apr 4 17:03:47 2022 AnchalSummaryCDSc1susaux2 slow controls acromag chassis fixed and installed

[Anchal, JC, Ian, Paco]

We have now fixed all issues with the PD mons of c1susaux2 chassis. The slow channels are now reading same values as the fast channels and there is no arbitrary offset. The binary channels are all working now except for LO2 UL which keeps showing ENABLE OFF. This was an issue earlier on LO1 UR and it magically disappeared and now is on LO2. I think the optical isolators aren't very robust. But anyways, now our watchdog system is fully functional for all BHD suspended optics.

Attachment 1: Screenshot_2022-04-04_17-03-26.png
Screenshot_2022-04-04_17-03-26.png
  16755   Mon Apr 4 15:49:06 2022 TegaUpdateOptical LeversSimplified sketch on BS table

I have updated the BS table using feedback from Koji and Paco and the attached pdf document is the latest iteration.

Attachment 1: BS_Table.drawio.pdf
BS_Table.drawio.pdf
  16754   Sat Apr 2 15:46:13 2022 ranaSummaryPEMParticle counter setup near BS Chamber

nice - please update the particle counter page in the 40m wiki. Its probably years out of date.

Quote:

I mounted the particle counter over the BS chamber attached to the cable tray as seen in Attachment 1. The signal cable runs through an active 30ft cable to the 1x2 rack. the wire is labeled and runs properly through the cable tray. The particle counter is plugged in at the power strip attached near the cable tray. The power cord is also labeled. 

 

  16753   Fri Apr 1 22:22:29 2022 KojiUpdateOptical LeversSimplified sketch on MC table

Possibility to swap BS and ITMX tables:
BS table, which Tega said MC table, is 2ft x 4ft. The ITMX table is 3ft x 5ft and only the central 2ft x 4ft area is used. The area around the BS table is the narrowest for the east arm. We need at least (2+delta) ft of the hallway width so that we can move the instrument. I'm not yet sure if the ITMX table can be placed there without precise investigation.
 

  16752   Fri Apr 1 17:02:02 2022 KojiUpdateOptical LeversSimplified sketch on MC table

We are supposed to have BS Oplev Beams. We don't like the shallow angle reflections (i.e. AOI>45deg).

The laser is too big but I suspect the other components are too small. So it'd be check the actual size of the components including the optical mounts that are missing on the figure so far.

  16751   Fri Apr 1 14:26:50 2022 TegaUpdateOptical LeversSimplified sketch on MC table

Here is an early sketch of the MC table. 

 

Quote:

I have made an editable draw.io diagram of the planned simplified BHD setup on the ITMY table (see attached).  10 pts = 1 inch.

This is very sketchy but easily adjustable since we are removing everything but the ITMY Oplev from that table.

 

Attachment 1: MC_Table.drawio.pdf
MC_Table.drawio.pdf
  16750   Fri Apr 1 14:26:19 2022 Ian MacMillanSummaryPEMParticle counter setup near BS Chamber

I mounted the particle counter over the BS chamber attached to the cable tray as seen in Attachment 1. The signal cable runs through an active 30ft cable to the 1x2 rack. the wire is labeled and runs properly through the cable tray. The particle counter is plugged in at the power strip attached near the cable tray. The power cord is also labeled. 

I restarted the particle counter service in the c1psl computer in the /etc/systemd/system/ folder using the commands

sudo systemctl restart particleCounter
sudo systemctl status particleCounter

I cannged the usb hub assigned in the service file to ttyUSB0 which is what we saw the computer had named it.

Checking the channels from this elog show the same particle count as when testing with the buttons and checking the screen. It seems that the channels had been down but are now restarted.

Attachment 1: IMG_1407.jpg
IMG_1407.jpg
  16749   Thu Mar 31 18:58:16 2022 Ian MacMillanSummaryBHDPart IIa of BHR upgrade - IR laser alignment on Xarm

[Ian, Paco]

We continued to try and lock the IR laser to the x-arm. Last time we adjusted TT2 PR2 and PR3 to adjust the beam to hit the center of the ITMX. Today we continued to adjust the same optics to make the beam hit the center of the beam splitter and make sure the beam is going in the direction of the ITMY. Then removed a mirror that wasn't even bolted down. I put it on the front right corner of the hood with all the other optics (right in front of a suspended mirror). With this unknown mirror removed the path was clear and hit the center of ITMX. From there we went and opened the ETMX chamber and using the IR scope we adjusted ITMX to put the beam on the center of the ETMX. We continued adjusting ETMX and ITMX until we had three reflections. We stopped here because we realized we should set up POX. We think this because any adjustment to set up the POX would mess up the work we had done on the x arm alignment.

  16748   Tue Mar 29 17:35:54 2022 PacoUpdateSUSDamping fix on BS, AS4, PR2, and PR3

[Ian, Paco]

  • We removed the "cheby" filters from AS4, PR2 and PR3 which had been misplaced after copying from the old SUS models. After removing them, the new SOS damped fine. Note that because of the Input matrices, the filters have to be enabled all at once for the MIMO loop to make sense.
  • We also disabled the "Cheby" filter on BS and saw it damp better. We don't understand this yet, but perhaps it's just a consequence of the many changes in the BSC that have rendered this filter obsolete.
  • we also reduced the damping gains on PR2, PR3 and AS4 to prevent overflow values. After the adjustments the optics were damping fine.
  16747   Tue Mar 29 15:07:11 2022 TegaSummaryVACOpening of MC and ETM chambers

[Anchal, Chub, Ian, Paco, Tega]

Today, we opened the MC chamber, ETMX chamber and ETMY chamber.

  16746   Mon Mar 28 16:25:34 2022 Ian MacMillanUpdateSUSETMY SUS Electronics Replacement - Questions

The point of changing the gains was to return the system to its origional state. ie I wanted the over all gain of the physical components to be the same as when we started. From the CDS side of things nothing else should be changed. The damping filters should remain at their origional values. The cts2um filter was changed to counteract a change in the electronics (replacing them). These changes should cancel eachother out. As for the side control, on 3/4/22 koji reduced the output resistors for the 4 face OSEMs but did not change the the SD one. there fore the SD did not need the same adjustment as the others.

Quote:

After Ian updated the cts2um filters for OSEM, shouldn't the damping gains be increased back by factor of 10 to previos values? Was the damping gain for SIDE ever changed? we found it at 250.

Can you explain why gain_offset filter was required and why this wasn't done for the side coil?

Quote:

I updated the gain of the ct2um filter on the OSEMS for ETMY and decreased their gain by a factor of 9 from 0.36 to 0.04.

I added a filter called "gain_offset" to all the coils except for side and added a gain of 0.48.

together these should negate the added gain from the electronics replacement of the ETMY

 

 

  16745   Sat Mar 26 13:04:38 2022 PacoSummaryBHDPart IIa of BHR upgrade - Green laser alignment on Yarm

I came for a bit to check on the "1Y0 burning smell" reported by Koji last evening. I agree that there's a toasty smell around the 1Y0 rack, but it is hard to tell where it comes from exactly. I think we can use the FLIR heat camera, but this will have to wait for Monday when someone with an iphone is around.

I also played a bit with the YAUX alignment (which is still poor) but I didn't try anything different nor did I seem to have more success. This is also an outstanding item for Monday, where maybe the next step is to install the GTRY path for a CCD.

  16744   Fri Mar 25 18:20:28 2022 AnchalSummaryBHDPart IIa of BHR upgrade - Green laser alignment on Yarm

[Anchal, Paco, Shruti]

  • Today we found out that the laser controller at the Yend had it's inputs disconnected for the FAST PZT input and the Slow temperature control input. The cables were also not labeled that were lying nearby.
  • I tracked the cables and connected the Fast input adn the slow input.
  • We then adjusted the temperature slider to go to previously marked laser crystal temperature of 40.1 degrees celsius. But we could not find the the beatnote of Y end IR with teh Main laser IR.
  • We scanned the laser crystal temeprature until we could see a beatntoe with the main IR laser and stopped in that region.
  • Then we continued looking for flashing and HOM, but no luck.
  • It was interesting that when the PDH loop was closed, we saw fluctuations in relfection photodiode DC value, but not so much when the loop was not closed.
  • We tried flipping the sign of the loop with no luck and we tried changing the gain of the loop as well, but no luck.
  • The behavior of this Yend laser and its controller is still suspicious to us. When the loop is closed, the PZT feedback signal should not cause a change in amplitude of the light so much as we see it.
  • Also the fact that when loop is clsoed we do not see signs of flashing in the reflection photodiode is interesting. If anyone has any ideas, that would be good.
  16743   Fri Mar 25 11:39:14 2022 AnchalUpdateSUSETMY SUS Electronics Replacement - Questions

After Ian updated the cts2um filters for OSEM, shouldn't the damping gains be increased back by factor of 10 to previos values? Was the damping gain for SIDE ever changed? we found it at 250.

Can you explain why gain_offset filter was required and why this wasn't done for the side coil?

Quote:

I updated the gain of the ct2um filter on the OSEMS for ETMY and decreased their gain by a factor of 9 from 0.36 to 0.04.

I added a filter called "gain_offset" to all the coils except for side and added a gain of 0.48.

together these should negate the added gain from the electronics replacement of the ETMY

 

  16742   Thu Mar 24 19:20:28 2022 AnchalSummaryBHDPart IIa of BHR upgrade - Green laser alignment on Yarm

[Anchal, Paco]


Uneloged: yesterday, Paco and Ian tried locking the green laser on Yend to the Y-arm. Nothing noteworthy happened, no luck.


  • Started with aligning the green injection to the Iris on table (it was misaligned) and getting it on the ITMY center.
  • Then we realized that the beam is going through the focusing leans very offceter. This makes the beam warped and deflected inside that needs to be corrected by mirrors.
  • So we decided to adjust the lens position first.
  • We realized that lens position is hard to adjust along with keeping the beam hiting ITMY.
  • So, we decided to remove the lens first and just align the green beam directly with the Iris on the table and the ITMY center. The beam became as big as ITMY when it reached there without the lens in path. We aligned it to fall on the mirror.
  • Then we placed the lens back so that the beam goes through the center of the lens.
  • Then we aligned ETMY to get the reflection back to the reflection PD.
  • But then when we checked the beam at ITMY again, it was severely missing ITMY. The possible reason is that ETMY is a diverging lens for the beam as well and changing its position changed the beam direction in the beam tube.
  • We decided to do another iteration of lens positioning with the new ETMY position. So same steps, removed lens, aligned beam with iris and ITMY, placed back the lens, aligned ETMY again to get reflection back on the PD.
  • This time the beam was still reaching ITMY. So we were glad that our iterations converged in 2 repititions.
  • Then we aligned ITMY to get counter propagating beam.
  • At the end of this alignment:
    • We see dips in relfection photodiode counts from 12200 to 9500.
    • The ETMY oplev is partially out of range (only 1 of the quadrant seeing light) but that could be because we changed the lens position. So Oplev servo is off.
  • We confirmed that the PDH servo is properly connected and loop is on. But we see no HOM catching lock for YARM even after us doing about 10 min search with ITMY-ETMY alignments.
  • We are leaving things as it is. We'll continue trying to align the cavity tomorrow. Everyone is welcome to try as long as they restore the positions of the optics to a good state at the end.

 

 

  16741   Mon Mar 21 18:42:06 2022 AnchalSummaryCDSc1auxey1 slow controls issues

Another issue, Green Y Shutter can not be controlled with the EPICS controls right now. This needs to be investigated.

  16740   Mon Mar 21 18:24:07 2022 AnchalSummaryBHDPart IIa of BHR upgrade - Green laser alignment on Yarm

[Anchal, Tega]

We did the following alignment steps on Yarm today:

  • We aligned the green beam going into the ETMY chamber to the three irises on the path.
  • We aligned ETMY to get counter-propagating prompt reflection from ETMY.
  • We were able to see light in the reflection photodiode for the green PDH loop.
  • Then we went to ITMY chamber and checked where the beam is coming to. It was coming very low in height.
  • We walked the beam on the end to get it near the center of ITMY.
  • Then we changed the ETMY alignment again to get back the reflection beam on the reflection photodiode.
  • Then we went again to ITMY chamber and changed the ITMY alignment to get counter-propagating reflection from ITMY.

After this, we came back to the control room. One peculiar thing to note is that the C1:LS-Y_REFL_DC_OUT channel is inverted i.e. it is showing negative values for the reflection DC voltage. On this signal, we see a little bit higher-order mode flashing but it is not bright enough to be seen on the face camera of the suspension. We'll continue aligning the cavity using CDS now to get TEM00 mode.


After trying for a bit, we were able to get flash of about 2000 counts which is about 16% of the max value of 12200 counts. We adjusted the ITMY angular position using the IFO_ALIGN screen but used the OPTICAL_ALIGN offset screen to adjust the angular position of ETMY. The pitch and yaw values for ITMY are 1.0360 and -0.260 respectively whereas the pitch and yaw values for ETMY are 5664.0 and -4937.0 respectively. 

  16739   Mon Mar 21 18:00:05 2022 Ian MacMillanUpdateBHDXARM AUX alignment cont.

[Ian, Paco]

Continuing with the previous alignment that we stoped on friday, we re set up my heavily cleaned iPhone on FaceTime. The Phone alowed us to see the laser on the ITMX and center it on that optic.

  • We increased the trigger level for the watchdog on the optics.
  • Green beam was centered on th ITMX.
  • aligned the green laser: prompt reflection off of the ETMX to maximize the signal on the reflection PD
  • Went back to ITMX chamber, looked for the beam in the tube to see if the laser was hitting the beam tube off of the ITMX. From there we comunicated how to adjust the beam to move the beam towards the ETMX.
  • On the ETMX the reflection PD signal was measured and the ITMX was adjusted using CDS until ripples/"flashes" were seen in the reflection PD channel (C1:ALS-X_REFL_DC_OUT_DQ).
  • Once we saw flashing we tried to minimize the reflection pd signal using only ITMX alignment slider so that all the light was being held in the cavity.
  • Once we had this going the PDH lock engaged the higher order mode that we were seeing.
  • We removed the phone setup and closed up the ITMX chamber.
  • From here we moved to the control room and continude to adjust the ITMX and ETMX to see if we could lock to a lower order mode.
    • This proved unsuccessful so we will resume by trying to replicate the ASS action manually (with M1 - M2 PZT sliders and the REFL PD)
Attachment 1: PXL_20220322_015247138.jpg
PXL_20220322_015247138.jpg
  16738   Mon Mar 21 14:22:52 2022 AnchalUpdateSUSETMY Alignmnet offsets needed to be changed

I'm not sure why but the PIT and YAW offset values of +2725 and -2341 were not sufficient for the reflected OPLEV beam to reach the OPLEV QPD. I had to change the C1:SUS-ETMY_PIT_OFFSET to 5641 and C1:SUS-ETMY_YAW_OFFSET to -4820 to come back to center of the OPLEV QPD. We aligned the Oplevs to center before venting, so hopefully this is our desired ETMY position.

On another note, the issue of ETMY unable to damp was simple. The alignment offsets were on to begin with with values above 1000. This meant that whenever we enabled coil output, ETMY would necessarily get a kick. All I had to do was keep alignment offsets off before starting the damping and slowly increase the alignment offsets to desired position.

Quote:
 

- This modification allowed me to align the oplev spot to the center of the QPD. C1:SUS-ETMY_PIT_OFFSET and C1:SUS-ETMY_YAW_OFFSE are +2725 (8%FS) and -2341 (7%FS), respectively.

  16737   Fri Mar 18 19:10:51 2022 AnchalSummaryCDSc1auxey1 slow controls issues

I started the modbusIOC service on c1auxey1 and added PD variance channels for UR and SD as well.  There are unfortunately two issues here:

  • The enable monitors are reading NOT of what they should read. The optical isolator circuit might need to be changed.
  • ETMY is not damping now. This is strange and was seen in the use to other acromag chassis as well where AS4 and PR2 are unable to damp. This is weird since the acromag chassis are not part of the damping loop, maybe it is a coincidence. Next time we should check if we still have this issue when acromag chassis is disconnected from ETMY.

 

  16736   Fri Mar 18 18:39:13 2022 YehonathanSummaryCDSc1auxey1 slow controls acromag chassis installed, powered

{Yehonathan, Anchal}

We connected the c1auxey1 chassie to the different boxes (coil drivers, SAT amp, etc.) using DB9 cables and labeling them in the process. We ran out of 2.5 foot DB9 cables so we used 5 foot as a temporary solution.

The chassie was powered, but a two issues arised:

1. The Acromags didn't turn on.

2. When connecting the green laser shutter BNC cable, the power supply overloaded.

We took the chassie back to the bench. The wire that powers the Acromags was disconnected. We made a new longer wire and made sure it is not connected flimsily.

The issue with the BNC turned out to be a much deeper problem: The GND and EXC wires on the DIN rail connector were switched! Making the shield of the BNC to have high volatage compared to the shield of the green shutter causing current to overflow when the BNC was connected.

We switched back the EXC and GND wires. Not trusting the digital I/O tests that were done before due to this mistake we tested some of the I/Os using a spare coil driver. We tested both the inputs and the outputs and they all seemed to work.

Finally, we also noticed that the 2 RTS DB9s were wrongly female type so we switched them to males. We closed the lead on chassie and installed it back in the rack. We connected the cables and saw that the green shutter BNC cable was no longer shorting the power supply.

  16735   Fri Mar 18 17:15:19 2022 PacoUpdateBHDXARM AUX alignment cont.

[Paco, Ian]

First, we re-balanced the ITMX OSEMs so that they would damp at around half-a-shadow.


Then, we set up a clean camera inside ITMX chamber looking at the ITMX optic. Then, using the live feed we aligned the AUX beam from the ETMX station using M1 and M2. The camera was great to help us align the beam properly close to the ITMX center. It wasn't very long until we could see a green beam on the IR card, but we didn't really see any flashing, so this may just be the bare transmission away from XARM resonance (Attachment #1).

Ian checked the reflection from ITMX using the IR card with holes, and he pretty much only saw one beam spot, so we turned to look for a beam scattering on the vacuum tube but didn't really see anything. This could mean that we were hitting the ETMX again, or missing slightly, or missing completely. We tried scanning the ITMX pitch and yaw using the bias (alignment) sliders, and with the illuminators off, try seeing some scattered green beam on the ETMX. We can't really see anything yet, but we will keep trying. If there are any tips on our method, it would be great to know them.

Attachment 1: Screen_Shot_2022-03-18_at_4.09.38_PM.png
Screen_Shot_2022-03-18_at_4.09.38_PM.png
  16734   Thu Mar 17 19:12:44 2022 AnchalSummaryCDSc1auxey1 slow controls acromag chassis installed, not powered

[Anchal, Tega]

We installed c1auxey1 computer and the acromag chassis in 1Y4. The computer has been configured properly for nfs mounts to happen and we have initialized a git repo for /cvs/cds/caltech/target/c1auxey1 directory which stores all files for running modbusIOC service on this computer. We connected 18V power source but have not connected the 24V power yet  as we need to make a new connector for it. Going on what Koji recommended, we'll connect the 24V power input to 18 V strip as well as the acromags can run on that voltage too.

ELOG V3.1.3-